OpenStudy (anonymous):

Find the number of minutes that Allegra went over the time that the plan allows. Explain your answer. Answer: Allegra went over 84 minutes that the plan allows because $8.40 was the overcharge of minutes, and it is 0.10 per minute so when you divide it the sum is 84 minutes.

OpenStudy (anonymous):

Please Help @Callisto @Hero

hero (hero):

Information from the original question is missing. Can't help without it.

OpenStudy (anonymous):

oh soory, can you check these. my teacher said the 3rd is wrong

OpenStudy (anonymous):

1. Allegra has a cell-phone plan that charges $65 per month and $0.10 for every minute that she uses the phone beyond what her plan allows. One month, she was billed $73.40. (a) Define a variable for the unknown. M=Minutes Used (b) Write an equation to model the problem. 73.40 = 65 + 0.10m (73.40 is the sum of 65 “Price per month” plus 0.10 “Price per overuse of minutes multiplied by the unknown amount of minutes used” (c) Solve the equation. Show your work. 73.40 = 65 + 0.10m 8.40 = 0.10m 84 = m (d) Find the number of minutes that Allegra went over the time that the plan allows. Explain your answer. Answer: Allegra went over 84 minutes that the plan allows because $8.40 was the overcharge of minutes, and it is 0.10 per minute so when you divide it the sum is 84 minutes. @Hero

hero (hero):

$8.40 is a dollar amount, not an amount of minutes.

OpenStudy (anonymous):

i dont understand

hero (hero):

You stated: $8.40 was the overcharge of minutes

hero (hero):

But dollars and minutes are two different quantities.